Leslie​ Mosallam, who recently sold her​ Porsche, placed $8,000 in a savings account paying annual compound interest of 4 percent.

  1. Calculate the amount of money that will accumulate if Leslie leaves the money in the bank for 2, 6 & 16 years
  2. Suppose Leslie moves her money into an account that pays 6 percent or one that pays 8 percent. Rework part (a) using 6 percent and 8 percent
  3. What conclusions can you draw about the relationship between interest​ rates, time, and future sums from the calculations you just​ did?
Answer #1

a) Initial amount = 8000

Interest rate = 4% compounded annually

Amount of money accumulated after 2 years is 8000*1.04^{2} = 8652.8

Amount of money accumulated after 6 years is 8000*1.04^{6} = 10122.55

Amount of money accumulated after 16 years is 8000*1.04^{16} = 14983.85

b) Going by the above calculation, Below are the values if the interest rates are increased to 6% and 8% respectively.

Amount of money accumulated after 2 years are 8000*1.06^{2} and 8000*1.08^{2} = 8988.8 and 9331.2

Amount of money accumulated after 2 years are 8000*1.06^{6} and 8000*1.08^{6} = 11348.15 and 12694.99

Amount of money accumulated after 2 years are 8000*1.06^{16} and 8000*1.08^{16} = 20322.81 and 27407.54

c) The power of compound interest is you earn interest on interest. Also, the benefit is more if you keep it for more number of years.

Even if the interest rates are doubled from 4% to 8%, the maximum benefit you will get only if you keep your investments for more number of years.